How do I find the Ka?

1.00 M of acetic acid with a pH of 2.16

0.100 M of acetic acid with a pH of 2.68

0.0100 M of acetic acid with a pH of 3.25
Answer #1

Ka = [H+] [A-] / [HA]

The square brackets indicate concentration, so the volumes are already included along with the moles of ions and undisociated acid.

pH gives concentration in moles per liter. [H+] = 10^-pH
[A-] = [H+]
[HA] = original concentration less [H+]

1). [H+] = 10^-pH

=10^-2.16 = 6.91*10^-3 = [A-]

   [HA] =1.00M

Ka=(6.91*10^-3)^2 / 1.00 = 4.77*10^-5

2). [H+] = 10^-pH

=10^-2.68 = 2.08*10^-3

Ka=( 2.08*10^-3)^2 / 0.1

=4.36*10^-5

3). [H+] = 10^-pH

=10^-3.25 = 5.62*10^-4

Ka= (5.62*10^-4)^2 / 0.01

=3.16*10^-5